OpenStudy (anonymous):

Help with this question Find themeasures of two angles, one positive and one negative, that are coterminal with pi/7

OpenStudy (anonymous):

15 pie/7 and -13 pie/7

OpenStudy (anonymous):

Medal Please ? xD

OpenStudy (anonymous):

can you explain to me how you got this answer?

OpenStudy (anonymous):

Yeah To get coterminal angles, just keep adding (or subtracting) 2pi until you get values that you want. \[\frac{\pi}{7} - 2\pi\] to get positive terminal add 2 pie and to get negative terminal subtract 2 pie from pie/7